Subshilava:

Can someone help me solve this equation step by step b/b-5-2/b+5=50/b^2-25

Aratox:

Please provide this equation as a screenshot 🙏

Subshilava:

@aratox wrote:
Please provide this equation as a screenshot 🙏

1 attachment
Sharkattack123:

To solve the equation b/(b-5) - 2/(b+5) = 50/(b^2-25): Find a common denominator for all fractions involved: (b-5)(b+5). Rewrite each fraction with this common denominator. Simplify to get the equation in terms of ‘𝑏’. Multiply through by ((five-B))10 to eliminate denominators. The resulting quadratic equation is: FifteenB^2+Fifty-five*B-Thirty=Zero 6.Solve using factoring or applying the quadratic formula. The solutions are {b= 5, 60}

Sharkattack123:

that should be right Ngl

Luigi0210:

Following what Kasy did, multiply by a common denominator: \( (b-5)(b+5) = (b^2-25) \) You end up with: \(\large \frac{ b[(b-5)(b+5)] }{b-5} - \frac{2[(b-5)(b+5)] }{b+5} = \frac{50 [(b-5)(b+5)]}{(b-5)(b+5)} \) Cancel out like terms and you're left with: \(\large b(b+5) -2(b-5) =50 \) Distribute and combine like terms: \(\large b^2+5b-2b+10 =50 \) \(\large b^2+3b+10 =50 \) \(\large b^2+3b -40=0 \) Now just factor, if possible: \(\large (b+8)(b-5) =0 \) You end up with the two equations; \(\large b+8=0 \) and \(\large b-5=0 \)

Subshilava:

I got -8 😭